Find Equivalent Annual Worth for the following Cash flow if i=3%:

Use excel to solve this question

show formula


Year Cash flow, $ 0 0 1 2 3 4 5 6 7 own 550 $50 350 Step 1. Finding NPV 5110 $130 Step 2. Finding AT
0 1 2 3 4 5 6 7 1 $50 $50 $50 $70 $90 $110 $130
Answer #1

We can determine the net present worth first then convert it to the Annual Equivalent

\large NPW = 50(P/A,3\%,2) + [50(P/A,3\%,5)+20(P/G,3\%,5)]\times (P/F,3\%,2)

\large \implies NPW = 50\times 1.9134+ [50\times 4.5797+20\times 8.8887]\times 0.9425

\large \implies NPW = 95.67 + [228.98 + 177.77]\times 0.9425

\large \implies NPW = 95.67 + 406.76\times 0.9425

\large \implies NPW = 95.67 + 383.41

\large \implies NPW = \$ \ 479.08

Now calculate the annual equivalent

\large AE = NPW(A/P,3\%,7)= \$ \ 479.08\times 0.1605

\large \implies AE \approx \$ \ 76.90

Refer the attached picture below the screenshot of excel sheet.

fx =NPV( CW68, CX60:CX66 ) CV CW СХ CY CZ Cash Flow Year 1.03 59 1 50 48.54368932 60 2 50 47.12979546 61 50 45.75708297 62 4

Please contact if having any query will be obliged to you for your generous support. Your help mean a lot to me please help. Thank you.